### v1

@@ I Introduction
Another major challenge is handling Non-Independent and Identically Distributed (Non-IID) data across clients. In real-world scenarios, clients often possess data that are not representative of the overall data distribution, leading to inconsistent global model performance and reduced convergence rates. This heterogeneity can cause the global model to perform poorly on some clients’ local data, undermining the benefits of collaborative learning [24]. The variations in local data distributions can result in conflicting model updates from different clients, making it difficult for the global model to generalize well across all participants.
Existing defence methods against backdoor attacks employ various strategies but are often limited by strict assumptions regarding data heterogeneity. Specifically, they assume data are Independent and Identically Distributed (IID) across clients and the proportion of malicious clients is relatively low. Robust aggregation methods, such as RLR [17] and FoolsGold [8], adjust the aggregation process to mitigate the influence of malicious updates but may struggle against sophisticated attacks and are less effective under Non-IID data distributions. Root data validation approaches, like FLTrust [4] and Baffle [2], attempt to validate updates using trusted data but can be circumvented by stealthy attacks that mimic benign behaviour. Differential Privacy methods, including FLAME [16], introduce noise and apply weight clipping to protect privacy and integrity but require precise tuning and often involve trade-offs that degrade model performance. Knowledge distillation-based methods, such as FedDF [13] and FedBE [5], address issues of Non-IID data and noisy gradients but may lack mechanisms for outlier detection, potentially allowing backdoor triggers to be transferred to the global model.
To overcome these limitations, we propose Robust Knowledge Distillation (RKD), a novel defense mechanism that enhances model integrity without relying on restrictive assumptions. RKD integrates clustering and model selection techniques to identify and filter out malicious updates, forming a reliable ensemble of models for knowledge distillation. Specifically, RKD employs cosine similarity and Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) to cluster client models and isolate potential outliers. Then, it selects models near the median of the benign cluster further to mitigate the influence of any remaining subtle malicious updates, thereby forming a trustworthy ensemble. Finally, it employs knowledge distillation to integrate the information from the ensemble into the global model, primarily reflecting benign update behaviour and mitigating subtle backdoor attacks, while also addressing Non-IID settings.
@@
  • We introduce the RKD framework, a robust aggregation method for FL that counters sophisticated backdoor attacks without relying on restrictive assumptions about data distribution or the number of malicious clients.
  • We conduct extensive evaluations of RKD on diverse datasets, including CIFAR-10, EMNIST, and Fashion-MNIST, demonstrating that RKD effectively mitigates various backdoor attacks such as Adversarially Adaptive Backdoor Attacks to Federated Learning (A3FL) [22], Focused-Flip Federated Backdoor Attack (F3BA) [7], Distributed Backdoor Attack (DBA) [20], and Train and Scale Backdoor Attack (TSBA) [3].
  • Our empirical analysis shows that RKD outperforms baseline defense methods, particularly in scenarios with high poisoning rates and varying levels of Non-IID data. RKD maintains high model accuracy above $80\%$ while reducing attack success rates to below $17\%$, underscoring its superior capability in maintaining FL resilience.
Extensive experiments underscore the effectiveness of RKD in enhancing FL robustness, and maintaining high model performance while effectively mitigating backdoor threats. The implementation of the RKD framework is available at https://github.com/AnonymeSc/RKD.
@@ II-C Characteristics of Backdoor Attacks
  • Angular Deviation: Malicious model updates may have a different direction in the parameter space compared to updates from benign clients [3]. This directional difference can be quantified using the cosine similarity (or angular deviation) between the parameter vectors of malicious updates $\theta_{\text{attack}}$ and benign updates $\theta_{\text{benign}}$: $\Delta\theta_{\text{angular}}=\cos^{-1}\left(\frac{\theta_{\text{attack}}\cdot\theta_{\text{benign}}}{\|\theta_{\text{attack}}\|\|\theta_{\text{benign}}\|}\right)$. Angular deviations indicate that the malicious update is contributing in a different direction, which may be due to the attacker’s attempt to embed the backdoor like F3BA attack [7].
@@ III Related Work
Clustering-Based Defences have been employed to distinguish between benign and malicious updates by grouping similar updates and identifying outlier models. Methods like FLAME [16] leverage HDBSCAN for clustering and incorporate noise addition to bolster security, yet they require improved adaptability to evolving threats. RFCL [1] also uses HDBSCAN to cluster clients and defend against gradient poisoning attacks but is less effective in defending against adaptive backdoor attacks [22]. A significant challenge with clustering-based defences is their reliance on the precision of the clustering process, which can be compromised in high-dimensional parameter spaces typical of deep learning models. Traditional clustering algorithms often struggle with high-dimensional data due to the curse of dimensionality, leading to less efficient clustering results and potential misclassification of benign updates as malicious or vice versa.
@@
RKD builds upon existing methods by innovatively integrating and enhancing foundational techniques such as HDBSCAN and knowledge distillation. This approach addresses the limitations of prior works, particularly in defending against recent backdoor attacks in Non-IID data distributions.
@@ IV-A Framework Overview
The RKD framework employs a multi-tiered strategy to enhance model robustness within FL. This subsequent provides an overview of the RKD framework, outlining its key components and processes. Detailed explanations of each strategy method are presented in the subsequent subsections.
Initially, the central server initializes the global model $M_{\text{global}}^{(0)}$ and broadcasts it to all participating clients. Each client $i$ trains its local model $M_{i}^{(r)}$ on its private dataset $D_{i}$, starting from the current global model $M_{\text{global}}^{(r)}$. The result is updated model parameters $\theta_{i}^{(r)}$, which are then returned to the server.
At the server, the focus is on identifying potential malicious updates. The server computes the cosine similarity between each client model update $\theta_{i}^{(r)}$ and the current global model $M_{\text{global}}^{(r)}$. This metric captures the angular similarity between local updates and the global model, enabling the detection of updates that significantly deviate in direction—a characteristic of malicious behaviour known as Angular Deviation.
Using these similarity scores, the server employs the HDBSCAN algorithm to cluster the models based on their similarity to the global model. This clustering process distinguishes benign updates, which form dense clusters due to their similarity, from potentially malicious updates, which appear as outliers because of their dissimilarity. By clustering based on similarity scores rather than directly using high-dimensional parameter vectors, the computational complexity is significantly reduced, and the clustering becomes more scalable and efficient.
Within the benign cluster, the server computes the median of the model parameters by taking the median of each parameter across the models. This helps mitigate the impact of extreme values in the updates, addressing the Magnitude Deviation characteristic of malicious behaviour.
@@
The selected models are then aggregated to form an initial distilled model. To further refine this model and improve its resistance to subtle backdoor triggers, while enhancing generalization across Non-IID data distributions, Knowledge Distillation (KD) is applied. The ensemble of selected models guides the refinement of the distilled model $M_{\text{global}}^{(r+1)}$, ensuring it replicates the collective behaviour of the ensemble models. This process helps protect against Subtle Deviations, where attackers craft updates that closely mimic benign updates in both magnitude and direction, making them harder to detect. Additionally, it improves performance across Non-IID.
Finally, the server broadcasts the refined global model $M_{\text{global}}^{(r+1)}$ to the identified benign clients for further training. By withholding the updated model from adversarial clients, RKD prevents attackers from adapting their strategies based on the latest changes, thereby reducing their ability to exploit subtle deviations. Clients classified as malicious continue to train using the previous global model $M_{\text{global}}^{(r)}$.
To further enhance security and mitigate the risk of adaptive attacks, RKD proposes an alternative solution. Instead of withholding the model entirely, the server sends a Perturbed Global Model (PGM) to clients suspected of being malicious This approach reduces the risk of adaptive attacks by obscuring the exact state of the global model while maintaining inclusivity, preventing attackers from deducing their classification. Malicious clients thus continue to train using this perturbed model, limiting their ability to exploit subtle deviations without providing them with precise information about the global model’s updates.
The perturbation added to the global model is intentionally kept minimal (e.g., noise scale of $1\times 10^{-4}$). This ensures that even if a benign client is mistakenly classified as malicious due to the non-IID nature of its data, the divergence from the true global model remains limited. The methodology allows for clients to be reclassified as benign in subsequent rounds. Once reclassified, they receive the unperturbed, updated global model, helping them realign with the overall training process.
Re-including previously misclassified benign clients may introduce slight variations in the global model’s performance. However, by keeping the perturbation scale minimal and allowing for dynamic reclassification, RKD mitigates this issue. The potential minor noise introduced is a reasonable trade-off for enhanced security against adaptive attacks.
@@
0:  Clients $\mathcal{A}$, number of iterations $R$
0:  Final global model $M_{\text{global}}^{(R)}$
1:  Initialize global model $M_{\text{global}}^{(0)}$
@@
3:     if  $r==0$  then
4:        Send $M_{\text{global}}^{(0)}$ to all clients in $\mathcal{A}$
@@
6:        Send $M_{\text{global}}^{(r)}$ to benign clients $\mathcal{A}_{\text{benign}}^{(r-1)}$
7:     end if
8:     Collect models ${M}^{(r)}=\{M_{i}^{(r)}\mid i\in\mathcal{A}\}$
9:     Identify ${M}_{\text{benign}}^{(r)}$ and $\mathcal{A}_{\text{benign}}^{(r)}$ via clustering $\triangleright$ See Algorithm 2
10:     Select ensemble models $\mathcal{E}^{(r)}$ from ${M}_{\text{benign}}^{(r)}$
11:     Compute aggregated model $M_{\text{distill}}^{(r)}$ from $\mathcal{E}^{(r)}$
12:     Update $M_{\text{global}}^{(r+1)}=\text{KD}(M_{\text{distill}}^{(r)},\mathcal{E}^{(r)})$triangleright See Algorithm 3
13:  end for
14:  return  Final global model $M_{\text{global}}^{(R)}$
Automated Clustering. This component identifies and excludes potentially malicious models to maintain the integrity of the federated learning process. The method leverages cosine similarity and the HDBSCAN clustering algorithm to distinguish between benign and malicious model updates.
At iteration $r$, each client $i$ has local model parameters $\theta_{i}^{(r)}$. To identify alignment with the current global model, the server computes the cosine similarity $s_{i}$ between each client’s model and the global model $M_{\text{global}}^{(r)}$, which serves as the reference point: $s_{i}=\frac{\left(\theta_{i}^{(r)}\right)^{\top}M_{\text{global}}^{(r)}}{\left\|\theta_{i}^{(r)}\right\|\left\|M_{\text{global}}^{(r)}\right\|},\quad\text{for }i=1,\ldots,N.$
Here, $M_{\text{global}}^{(r)}$ represents the current aggregated global model. Higher cosine similarity values indicate greater alignment with the global model, which is expected for benign clients, while malicious clients are more likely to deviate.
@@
This dynamic adjustment ensures that the clustering remains adaptable to changes in client participation and training dynamics. For example, as the training progresses and fewer clients actively participate, $Q$ decreases, maintaining meaningful cluster sizes and computational efficiency.
Applying HDBSCAN to the set of similarity scores yields cluster labels $\{L_{i}\}$: $\{L_{i}\}=\text{HDBSCAN}\left(\{s_{i}\},Q\right).$
For each identified cluster $C_{k}$, the mean cosine similarity $\mu_{k}$ is computed: $\mu_{k}=\frac{1}{\left\|C_{k}\right\|}\sum_{i\in C_{k}}s_{i},$ where $\left\|C_{k}\right\|$ is the number of clients in cluster $C_{k}$.
@@
Cluster $\text{Cluster }C_{k}\text{ is classified as }\begin{cases}{M}_{\text{benign}},&\text{if }\mu_{k}=\mu_{\text{max}},\\ {M}_{\text{malicious}},&\text{otherwise}.\end{cases}$
Models in the benign cluster are used for further training and receive the updated global model in the next iteration. Models in malicious clusters are excluded from receiving the updated global model, limiting their influence on the federated learning process. The detailed steps of the automated clustering procedure are in Algorithm 2.
@@
0:  ${M}^{(r)}=\{\theta_{i}^{(r)}\}_{i=1}^{N}$: Client models at iteration $r$
0:  ${M}_{\text{benign}}^{(r)}$, $\mathcal{A}_{\text{benign}}^{(r)}$: Benign models and client indices
@@
2:     Compute cosine similarity score $s_{i}=\frac{\left(\theta_{i}^{(r)}\right)^{\top}M_{\text{global}}^{(r)}}{\left\|\theta_{i}^{(r)}\right\|\left\|M_{\text{global}}^{(r)}\right\|}$
@@
8:  Determine benign cluster $\mu_{\text{max}}=\max\left(\{\mu_{k}\}\right)$
9:  Identify benign clients $\mathcal{A}_{\text{benign}}^{(r)}=\left\{i\mid L_{i}=k\text{ and }\mu_{k}=\mu_{\text{max}}\right\}$
10:  Collect benign models ${M}_{\text{benign}}^{(r)}=\left\{\theta_{i}^{(r)}\mid i\in\mathcal{A}_{\text{benign}}^{(r)}\right\}$
11:  return  ${M}_{\text{benign}}^{(r)}$, $\mathcal{A}_{\text{benign}}^{(r)}$
Model Selection. This component refines the set of benign models to select the most representative ones for aggregation. By focusing on models closest to the central tendency, it mitigates the impact of outliers and enhances the robustness of the aggregated model against backdoor attacks.
From the set of benign models ${M}_{\text{benign}}^{(r)}=\left\{\theta_{i}^{(r)}\mid i\in\mathcal{A}_{\text{benign}}^{(r)}\right\}$, we compute the median model parameter vector $\theta_{\text{median}}^{(r)}$. This is done by taking the median of each parameter across all benign models: $\theta_{\text{median}}^{(r)}=\text{median}\left(\left\{\theta_{i}^{(r)}\mid i\in\mathcal{A}_{\text{benign}}^{(r)}\right\}\right),$ where the median is computed element-wise over the parameter vectors.
Next, we calculate the distance between each benign model $\theta_{i}^{(r)}$ and the median model $\theta_{\text{median}}^{(r)}$ using the $L_{1}$ norm: $d_{i}=\left\|\theta_{i}^{(r)}-\theta_{\text{median}}^{(r)}\right\|_{1},\quad\text{for all }i\in\mathcal{A}_{\text{benign}}^{(r)}.$
Models are then ranked based on their distances $d_{i}$. A subset of models with the smallest distances is selected to form the ensemble $\mathcal{E}^{(r)}$: $\mathcal{E}^{(r)}=\left\{\theta_{i}^{(r)}\in{M}_{\text{benign}}^{(r)}\mid d_{i}\leq\epsilon\right\},$ where $\epsilon$ is a statistical measure of the standard deviation of distances from the median to inform the setting of epsilon, ensuring it adapts to the inherent variability in benign updates.
Knowledge Distillation Process. It refines the global model by distilling knowledge from an ensemble of selected benign models, denoted as $\mathcal{E}^{(r)}$, which have been identified through automated clustering and model selection in previous steps. The server utilizes an unlabeled dataset $D_{\text{val}}$, comprising $16\%$ of the total training data, to perform knowledge distillation. Pseudo-labels are generated from the ensemble model outputs to guide the training of the distilled model $M_{\text{distill}}$.
For each sample $x$ in $D_{\text{val}}$, the server computes the logits using each model in the ensemble. Here, logits refer to the raw, unnormalized output scores from the models before applying the softmax function. These logits are averaged to produce the ensemble logits:
@@
where $f_{M_{i}}(x)$ denotes the logits produced by model $M_{i}$ for input $x$. The pseudo-labels are then generated by applying the softmax function with a temperature parameter $T$ to the averaged logits:
@@
The temperature parameter $T$ in the softmax function serves as a scaling factor that controls the smoothness of the resulting probability distribution.
The distilled model $M_{\text{distill}}$ is then trained to minimize the Kullback-Leibler (KL) divergence between its output probabilities and the pseudo-labels:
@@
Here, $f_{M_{\text{distill}}}(x)$ represents the logits from the distilled model for input $x$. The KL divergence measures the difference between the two probability distributions, encouraging the distilled model to align closely with the ensemble’s behaviour.
To further stabilize and enhance the generalization of $M_{\text{distill}}$, we employ Stochastic Weight Averaging (SWA) during training. SWA improves generalization by maintaining a running average of the model weights over training epochs, effectively capturing the trajectory of the model parameters as they converge.
During each epoch of the knowledge distillation training, after updating the distilled model $M_{\text{distill}}$ using gradient descent, the SWA model $M_{\text{SWA}}$ is updated as follows:
@@
where $n_{\text{SWA}}$ is the number of times the SWA model has been updated. Initially, $M_{\text{SWA}}$ is set to the weights of $M_{\text{distill}}$, and $n_{\text{SWA}}=1$.
Finally, the SWA model $M_{\text{SWA}}$ is set as the updated global model for the next iteration: $M_{\text{global}}^{(r+1)}\leftarrow M_{\text{SWA}}.$
This updated global model integrates the distilled knowledge from the ensemble and the generalization properties provided by SWA. The process of knowledge distillation is in Algorithm 3.
@@
0:  $\mathcal{E}^{(r)}$: Ensemble of selected benign models; $D_{\text{val}}$: Unlabeled data for distillation; $T$: Temperature for softmax; $E_{\text{KD}}$: Number of epochs for knowledge distillation; $\eta$: Learning rate
0:  Updated global model $M_{\text{global}}^{(r+1)}$
@@
2:  Generate pseudo-labels: $\tilde{y}(x)=\text{softmax}\left(\frac{\text{Ensemble\_Logits}(x)}{T}\right)$
3:  for epoch $e=1$ to $E_{\text{KD}}$ do
4:     for each mini-batch $\{x_{b}\}\subset D_{\text{val}}$ do

### v2

@@ I Introduction
Another significant challenge arises from the Non-IID data distribution across clients. In real-world scenarios, client data often differ substantially from the overall data distribution, leading to inconsistent global model performance and slower convergence [24]. This heterogeneity can cause conflicting model updates from different clients, making it difficult for the global model to generalize well across all participants. Many existing defense methods assume that client data are Independent and Identically Distributed (IID) and that only a small fraction of clients are malicious [27],[28],[8]. These assumptions simplify the aggregation process but fail to capture the complexity of practical FL deployments.
Various defense strategies have been explored. Robust aggregation methods, such as RLR [17] and FoolsGold [8], modify the aggregation process to reduce the influence of malicious updates; however, they may struggle against sophisticated backdoor attacks and perform suboptimally in Non-IID settings. Other approaches include root data validation methods like FLTrust [4] and Baffle [2], which use trusted data to validate client updates, yet they can be circumvented by adversaries that mimic benign behavior. Differential Privacy-based methods, exemplified by FLAME [16], incorporate noise and weight clipping to maintain privacy and integrity, but they require careful tuning and often compromise model performance. While approaches such as FedDF [13] and FedBE [5] have been proposed to address the challenges posed by Non-IID data—primarily through knowledge distillation techniques—they focus on improving model performance under data heterogeneity rather than explicitly defending against backdoor attacks. In other words, although these methods help mitigate issues caused by Non-IID distributions, they do not provide targeted defenses against adversarial updates.
To bridge this gap, we propose Robust Knowledge Distillation (RKD), a novel defense mechanism that enhances the integrity of the global model without relying on restrictive IID assumptions or low malicious-client ratios. RKD integrates clustering and model selection techniques to identify and filter out malicious updates, forming a reliable ensemble of benign models. Specifically, it employs cosine similarity measures and Hierarchical Density-Based Spatial Clustering of Applications with Noise (HDBSCAN) to cluster client models, isolating potential outliers. Models closest to the median of the benign cluster are then selected to form an ensemble that is subsequently distilled into the global model. This approach defends against backdoor attacks in Non-IID settings.
@@
  • We introduce the RKD framework, which robustly aggregates client updates by filtering out adversarial contributions without assuming IID data or a low fraction of malicious clients.
  • We evaluate RKD on diverse datasets, including CIFAR-10, EMNIST, and Fashion-MNIST, demonstrating its effectiveness against sophisticated backdoor attacks such as A3FL [22], F3BA [7], DBA [20], ADBA [26], and TSBA [3].
  • Empirical results show that RKD maintains high model accuracy above 80% while reducing attack success rates below 17%, outperforming state-of-the-art defense methods, particularly in challenging Non-IID scenarios.
Extensive experiments underscore the effectiveness of RKD in enhancing FL robustness, and maintaining high model performance while effectively mitigating backdoor threats. The implementation of the RKD framework is available at https://github.com/EbtisaamCS/RKD.git.
@@ II-C Characteristics of Backdoor Attacks
  • Angular Deviation: Malicious model updates may have a different direction in the parameter space compared to updates from benign clients [3]. This directional difference can be quantified using the cosine similarity (or angular deviation) between the parameter vectors of malicious updates $\theta_{\text{attack}}$ and benign updates $\theta_{\text{benign}}$: $\Delta\theta_{\text{angular}}=\cos^{-1}\left(\frac{\theta_{\text{attack}}\cdot\theta_{\text{benign}}}{\|\theta_{\text{attack}}\|\|\theta_{\text{benign}}\|}\right)$. Angular deviations indicate that the malicious update is contributing in a different direction, which may be due to the attacker’s attempt to embed the backdoor, such as F3BA attack [7].
@@ III Related Work
Clustering-Based Defences. Prior work has explored clustering-based techniques to distinguish between benign and malicious updates in federated learning by grouping similar updates and flagging outliers. For example, FLAME [16] leverages HDBSCAN for clustering and introduces noise to enhance security; however, its adaptability to evolving threat landscapes remains limited. Similarly, RFCL [1] employs HDBSCAN to cluster client updates for defending against gradient poisoning attacks, yet its effectiveness diminishes against adaptive backdoor attacks [22]. A critical limitation of these methods lies in their reliance on the precision of the clustering process, which is often compromised in high-dimensional parameter spaces typical of deep learning models. In such spaces, traditional clustering algorithms can struggle due to the curse of dimensionality, leading to inefficient clustering and the potential misclassification of benign updates as malicious, or vice versa.
In contrast, our approach mitigates these issues by first computing cosine similarity scores between client updates and the global model. These scalar similarity values capture the directional alignment of updates while reducing the data from a high-dimensional space to a one-dimensional representation, thereby simplifying the clustering task. We then apply HDBSCAN to these cosine similarity scores, which significantly enhances clustering efficiency and accuracy.
@@
Our proposed approach, RKD, builds upon existing methods by innovatively integrating and enhancing foundational techniques such as HDBSCAN and knowledge distillation. This approach addresses the limitations of prior works, particularly in defending against recent backdoor attacks in Non-IID data.
@@ IV-A Framework Overview
The proposed RKD framework employs a multi-tiered strategy to enhance the robustness of federated learning systems. We provide an overview of the framework by outlining its key components and processes. Detailed explanations of each component are presented in the subsequent subsections.
Initially, the central server initializes the global model $M_{\text{global}}^{0}$ and broadcasts it to all participating clients. Each client $i$ trains its local model $M_{i}^{r}$ on its private dataset $D_{i}$, starting from the current global model $M_{\text{global}}^{r}$. The resulting updated model parameters $\theta_{i}^{r}$ are then returned to the server.
At the server, the focus is on identifying potential malicious updates. The server computes the cosine similarity between each client model update $\theta_{i}^{r}$ and the current global model $M_{\text{global}}^{r}$. This metric captures the angular similarity between local updates and the global model, enabling the detection of updates that significantly deviate in direction—a characteristic of malicious behavior known as Angular Deviation.
Using these similarity scores, the server employs the HDBSCAN algorithm to cluster the models based on their similarity to the global model. This clustering process distinguishes benign updates, which form dense clusters due to their similarity, from potentially malicious updates, which appear as outliers because of their dissimilarity. By clustering based on similarity scores rather than directly using high-dimensional parameter vectors, the computational complexity is significantly reduced, making the clustering process more scalable and efficient.
Within the benign cluster, the server computes the median of the model parameters by taking the median of each parameter across the models. This helps mitigate the impact of extreme values in the updates, addressing the Magnitude Deviation characteristic of malicious behavior.
@@
The selected models are then aggregated to form an initial distilled model. To further refine this model and improve its resilience against subtle backdoor triggers in Non-IID settings, Knowledge Distillation (KD) is applied. The ensemble of selected models guides the refinement of the distilled model $M_{\text{global}}^{r+1}$, ensuring it reflects the collective behavior of benign clients. This process mitigates the risk of Subtle Deviations, where attackers craft updates that closely mimic benign updates in both magnitude and direction.
Finally, the server broadcasts the refined global model $M_{\text{global}}^{r+1}$ to benign clients for further training. For clients identified as malicious, our primary approach is the Exclusion Strategy: these clients are not provided with the updated global model $M_{\text{global}}^{r+1}$; instead, they continue training using their current local model until they are reclassified as benign. This complete exclusion prevents adversaries from adapting their strategies based on the most recent global model updates.
As an alternative, we also consider the Perturbation Strategy (referred to as RKD Perturbed Global Model, or PGM). In this approach, to mask the fact that a client has been flagged as malicious, the server sends a perturbed version of the refined global model. Specifically, the perturbed model is defined as $M_{\text{pert}}^{r+1}=M_{\text{global}}^{r+1}+\eta,$ where $\eta$ is a noise vector with a small magnitude (e.g., $\|\eta\|\approx 1\times 10^{-4}$). This strategy limits the opportunity for adaptive adversaries to deduce their status while still providing them with an updated model.
In both cases, if a client previously classified as malicious is reclassified as benign in subsequent rounds, it will begin receiving the updated global model $M_{\text{global}}^{r+1}$ to realign with the overall training process.
@@
[…]
0:  Final global model $M_{\text{global}}^{R}$
1:  Initialize global model $M_{\text{global}}^{0}$
@@
3:     if $r=0$ then
4:        Send $M_{\text{global}}^{0}$ to all clients in $\mathcal{A}$
@@
6:        Send $M_{\text{global}}^{r}$ to benign clients $\mathcal{A}_{\text{benign}}^{r-1}$
7:        if $S=\text{Exclusion}$ then
8:           for each malicious client $i\in\mathcal{A}\setminus\mathcal{A}_{\text{benign}}^{r-1}$ do
9:              Send the current local model $M_{i}^{r}$ to client $i$
10:           end for
11:        else {Otherwise, using Perturbation}
12:           Compute perturbed model $M_{\text{pert}}^{r}=M_{\text{global}}^{r}+\eta$
13:           Send $M_{\text{pert}}^{r}$ to malicious clients $\mathcal{A}\setminus\mathcal{A}_{\text{benign}}^{r-1}$
14:        end if
15:     end if
16:     Collect models $\{M^{r}\}=\{M_{i}^{r}\mid i\in\mathcal{A}\}$
17:     Identify $\{M_{\text{benign}}^{r}\}$ and $\mathcal{A}_{\text{benign}}^{r}$ via clustering $\triangleright$ See Algorithm 2
18:     Select ensemble models $\mathcal{E}^{r}$ from $\{M_{\text{benign}}^{r}\}$
19:     Compute aggregated model $M_{\text{distill}}^{r}$ from $\mathcal{E}^{r}$
20:     Update $M_{\text{global}}^{r+1}=\text{KD}(M_{\text{distill}}^{r},\mathcal{E}^{r})$triangleright See Algorithm 3
21:  end for
22:  return  Final global model $M_{\text{global}}^{R}$
IV-B Automated Clustering.
This component identifies and excludes potentially malicious models to maintain the integrity of the federated learning process. The method leverages cosine similarity and the HDBSCAN clustering algorithm to distinguish between benign and malicious model updates.
At iteration $r$, each client $i$ has local model parameters $\theta_{i}^{r}$. To identify alignment with the current global model, the server computes the cosine similarity $s_{i}$ between each client’s model and the global model $M_{\text{global}}^{r}$:
$s_{i}=\frac{\left(\theta_{i}^{r}\right)^{\top}M_{\text{global}}^{r}}{\left\|\theta_{i}^{r}\right\|\left\|M_{\text{global}}^{r}\right\|},\quad\text{for }i=1,\ldots,N.$
Higher cosine similarity values indicate greater alignment with the global model, which is expected for benign clients, while malicious clients are more likely to deviate.
@@
Applying HDBSCAN to the set of similarity scores yields cluster labels $\{L_{i}\}$: $\{L_{i}\}=\text{HDBSCAN}\left(\{s_{i}\},Q\right).$ For each identified cluster $C_{k}$, the mean cosine similarity $\mu_{k}$ is computed: $\mu_{k}=\frac{1}{\left\|C_{k}\right\|}\sum_{i\in C_{k}}s_{i},$ where $\left\|C_{k}\right\|$ is the number of clients in cluster $C_{k}$.
@@
Cluster $\text{Cluster }C_{k}\text{ is classified as }\begin{cases}\text{benign},&\text{if }\mu_{k}=\mu_{\text{max}},\\ \text{malicious},&\text{otherwise}.\end{cases}$
Models in the benign cluster are used for further training and receive the updated global model in the next iteration. For models identified as malicious, our framework supports two strategies. In the primary Exclusion Strategy, malicious clients are entirely excluded from receiving the updated global model, thereby limiting their influence on the federated learning process. Alternatively, under the Perturbation Strategy, malicious clients receive a perturbed version of the updated global model, which similarly limits their influence while obscuring their classification.
The detailed steps of the automated clustering procedure are presented in Algorithm 2.
@@
0:  $\{M^{r}\}=\{\theta_{i}^{r}\}_{i=1}^{N}$: Client models at iteration $r$
0:  $\{M_{\text{benign}}^{r}\}$, $\mathcal{A}_{\text{benign}}^{r}$: Benign models and client indices
@@
2:     Compute cosine similarity score $s_{i}=\frac{\left(\theta_{i}^{r}\right)^{\top}M_{\text{global}}^{r}}{\left\|\theta_{i}^{r}\right\|\left\|M_{\text{global}}^{r}\right\|}$
@@
8:  Determine benign cluster: $\mu_{\text{max}}=\max\left(\{\mu_{k}\}\right)$
9:  Identify benign clients: $\mathcal{A}_{\text{benign}}^{r}=\left\{i\mid L_{i}=k\text{ and }\mu_{k}=\mu_{\text{max}}\right\}$
10:  Collect benign models: $\{M_{\text{benign}}^{r}\}=\left\{\theta_{i}^{r}\mid i\in\mathcal{A}_{\text{benign}}^{r}\right\}$
11:  return  $\{M_{\text{benign}}^{r}\}$, $\mathcal{A}_{\text{benign}}^{r}$
IV-C Model Selection.
This component refines the set of benign models by selecting the most representative ones for aggregation, thereby mitigating the impact of outliers and enhancing the robustness of the aggregated model against backdoor attacks.
From the set of benign models $\{M_{\text{benign}}^{r}\}=\{\theta_{i}^{r}\mid i\in\mathcal{A}_{\text{benign}}^{r}\},$ we compute the median model parameter vector $\theta_{\text{median}}^{r}$ by taking the element-wise median across all benign models: $\theta_{\text{median}}^{r}=\text{median}\left(\{\theta_{i}^{r}\mid i\in\mathcal{A}_{\text{benign}}^{r}\}\right).$
Next, we calculate the distance between each benign model $\theta_{i}^{r}$ and the median model $\theta_{\text{median}}^{r}$ using the $L_{1}$ norm: $d_{i}=\left\|\theta_{i}^{r}-\theta_{\text{median}}^{r}\right\|_{1},\quad\text{for all }i\in\mathcal{A}_{\text{benign}}^{r}.$
These distances $\{d_{i}\}$ are then used to rank the benign models. To select the most representative models, we define a threshold $\epsilon$ based on the dispersion of the $d_{i}$ values. For example, one may set $\epsilon=\mu_{d}+k\sigma_{d},$ where $\mu_{d}$ and $\sigma_{d}$ denote the mean and standard deviation of the distances $\{d_{i}\}$, and $k$ is a hyperparameter that controls the threshold level. The ensemble is then defined as: $\mathcal{E}^{r}=\left\{\theta_{i}^{r}\in\{M_{\text{benign}}^{r}\}\mid d_{i}\leq\epsilon\right\}.$ This threshold ensures that only models whose distances from the median are within a reasonable range are selected, thereby accommodating the inherent variability in benign updates.
IV-D Knowledge Distillation Process.
This process refines the global model by distilling knowledge from an ensemble of selected benign models $\mathcal{E}^{r}$, identified through the previous steps. The server utilizes an unlabeled dataset $D_{\text{val}}$ (which is separate from the clients’ training data and comprises 16% of the total training data) for knowledge distillation. Pseudo-labels are generated from the ensemble’s outputs to guide the training of the distilled model.
For each sample $x$ in $D_{\text{val}}$, the server computes the logits using each model in the ensemble. These logits are averaged to produce the ensemble logits:
@@
where $f_{M_{i}}(x)$ denotes the logits (i.e., the raw output scores before softmax) produced by model $M_{i}$ for input $x$. The pseudo-labels are then generated by applying the softmax function with a temperature parameter $T$:
@@
Dividing by $T$ smooths the probability distribution, allowing for a softer target during distillation.
The distilled model $M_{\text{distill}}$ is trained to minimize the Kullback-Leibler (KL) divergence between its output probabilities and the pseudo-labels:
@@
This loss encourages the distilled model to align with the behavior of the ensemble.
To further stabilize training and enhance generalization, we employ Stochastic Weight Averaging (SWA). SWA maintains a running average of the model weights, capturing the trajectory of the parameters as they converge. During each epoch of knowledge distillation, after updating $M_{\text{distill}}$ using gradient descent, the SWA model $M_{\text{SWA}}$ is updated as follows:
@@
where $n_{\text{SWA}}$ is the number of SWA updates. Initially, $M_{\text{SWA}}$ is set to $M_{\text{distill}}$ and $n_{\text{SWA}}=1$.
Finally, after completing the distillation process, the SWA model is set as the updated global model for the next iteration:
$M_{\text{global}}^{r+1}\leftarrow M_{\text{SWA}}.$
The process of knowledge distillation is summarized in Algorithm 3.
@@
0:  $\mathcal{E}^{r}$: Ensemble of selected benign models; $D_{\text{val}}$: Unlabeled data for distillation; $T$: Temperature for softmax; $E_{\text{KD}}$: Number of epochs; $\eta$: Learning rate
0:  Updated global model $M_{\text{global}}^{r+1}$
@@
2:  Generate pseudo-labels:
$\tilde{y}(x)=\text{softmax}\left(\frac{\text{Ensemble\_Logits}(x)}{T}\right)$
3:  Initialize $M_{\text{SWA}}\leftarrow M_{\text{distill}}$ and $n_{\text{SWA}}\leftarrow 1$
4:  for epoch $e=1$ to $E_{\text{KD}}$ do
5:     for each mini-batch $\{x_{b}\}\subset D_{\text{val}}$ do